# GIMLET: Generalizable and Interpretable Model Learning through Embedded Thermodynamics

### v1

@@ Abstract
We develop a data-driven framework for discovering constitutive relations in models of fluid flow and scalar transport. Our approach infers unknown closure terms in the governing equations (gray-box discovery) under the assumption that the temporal derivative, convective transport, and pressure-gradient contributions are known. The formulation is rooted in a variational principle from nonequilibrium thermodynamics, where the dynamics is defined by a free-energy functional and a dissipation functional. The unknown constitutive terms arise as functional derivatives of these functionals with respect to the state variables. To enable a flexible and structured model discovery, the free-energy and dissipation functionals are parameterized using neural networks, while their functional derivatives are obtained via automatic differentiation. This construction enforces thermodynamic consistency by design, ensuring monotonic decay of the total free energy and non-negative entropy production. The resulting method, termed GIMLET (Generalizable and Interpretable Model Learning through Embedded Thermodynamics), avoids reliance on a predefined library of candidate functions, unlike sparse regression or symbolic identification approaches. The learned models are generalizable in that functionals identified from one dataset can be transferred to distinct datasets governed by the same underlying equations. Moreover, the inferred free-energy and dissipation functions provide direct physical interpretability of the learned dynamics. The framework is demonstrated on several benchmark systems, including the viscous Burgers equation, the Kuramoto–Sivashinsky equation, and the incompressible Navier–Stokes equations for both Newtonian and non-Newtonian fluids.
@@ 1 Introduction
Neural operator style models aim to learn mappings between function spaces, such as the evolution of fields over time. Representative methods include Neural ODE [Chen2018], Hamiltonian Neural Networks [Greydanus2019], Fourier Neural Operator [Li2020], DeepONet [Lu2021], and many others [Lusch2018, Champion2019, Pfister2019, Yuan2019]. These approaches are largely black-box and often offer limited interpretability. In terms of generalizability, they typically require training on diverse data spanning many operating conditions, which can be time consuming.
In contrast, symbolic regression methods lie closer to the white box end of the spectrum and aim to recover explicit governing equations from data. Early successes in data-driven equation discovery were achieved using genetic programming-based symbolic regression, which represents candidate governing equations as expression trees and searches for parsimonious models consistent with data [Bongard2007, Schmidt2009]. Symbolic regression (SR) comprises methods that discover mathematical expressions that best fit a given dataset [schmidt2009symbolic, stoutemyer2013can]. Among SR methods, genetic programming is one of the most widely used techniques, in which expression trees are iteratively modified using elementary mathematical operators to produce concise, data-consistent models. It expresses candidate models as combinations of elementary operators, variables, and constants, yielding explicit analytical forms. In genetic programming, candidate expressions are represented as tree structures, where internal nodes correspond to operators and leaf nodes represent variables or constants. Model discovery maintains a population of such trees and iteratively improves them through selection, mutation, and crossover based on their agreement with data. In recent work by [cranmer2023interpretable], both the performance and computational efficiency of symbolic regression were further enhanced through an island-based evolutionary framework. In this approach, populations were partitioned into multiple islands, each characterized by distinct operator sets and mutation strategies. Genetic algorithm operations, including mutation and crossover, were performed both within individual islands and across islands, promoting population diversity, improved exploration of the search space, and more effective convergence toward interpretable symbolic models.
@@
As reviewed so far, existing methods for physical model discovery typically fall short in at least one of three aspects: generalizability, interpretability, or library-free. We propose a new methodology that aims to achieve all three objectives. To this end, we leverage thermodynamics and a variational principle. Thermodynamics is among the most universal and transferable descriptions of physical systems. Enforcing thermodynamic consistency can lead to more robust constitutive laws, stable numerical schemes, and physically meaningful discovered models [Cueto2023]. Motivated by this idea, several machine learning frameworks have been proposed, including Themodynamics-based Artificial Neural Networks (TANNs) [Masi2021] and thermodynamically consistent physics-informed neural networks [Patel2022]. Motivated by the same goal of enforcing thermodynamic consistency in a general and systematic way, the General Equation for Non-Equilibrium Reversible Irreversible Coupling (GENERIC) framework was developed as a universal description of macroscopic dynamics. In GENERIC, the evolution equations are constructed from energy and entropy potentials, providing a unified structure that couples reversible and irreversible processes and can be viewed as a thermodynamically grounded extension of Ginzburg-Landau equation [Grmela1997, Oettinger1997]. Based on the GENERIC concept, Zhang2022 proposed the GENERIC formalism-informed neural networks (GFINNs) for physical model discovery. However, the applicability of GFINN has so far been investigated primarily for ordinary differential equations, leaving extensions to PDE settings open.
According to the variational principle, also called the principle of least action path, realizable motions correspond to the stationary points of an action functional. Onsager’s variational principle [Onsager1931a, Onsager1931], which formulated the time evolution of the free energy and dissipation potentials, can lead to thermodynamically consistent governing equations for non-equilibrium dynamics [Van2020]. Building on this principle, Huang2022 proposed the Variational Onsager Neural Networks (VONNs), which aim to discover the free energy and dissipation potentials as neural networks. Although VONNs are a sophisticated and novel methodology, their application is limited to systems whose dynamics can be fully described by a free energy and dissipation potentials. This restriction makes direct application to fluid flow challenging. For instance, applying Onsager’s variational principle to the Newtonian fluid in a straightforward way yields only a linearized form of the Navier-Stokes equation [Takaki2007]. In addition, the VONNs framework assumes that both the state variables and process variables are available as training data, which is often unrealistic for practical experimental observations.
@@ 2 Methodology
This section presents our thermodynamically consistent, library-free framework for discovering missing terms in continuum PDEs.
2.1 Basic concept
We consider physical law discovery in continuum mechanics, where spatiotemporal field variables are evolving according to partial differential equations (PDEs). We assume that part of the governing equations is known and focus only on the missing terms (gray-box discovery). To represent these unknown contributions, an additional free energy function $g$ and dissipation function $\theta$ are introduced. Basic forms of governing equations (GEs) are derived through the variational principle. In the resulting GEs, the unknown terms are expressed as functional derivatives of $g$ and $\theta$. To learn $g$ and $\theta$ from data, we employ physics-informed neural networks (PINNs). Using PINNs, the functional derivatives for unknown functions can be calculated through automatic differentiation. The total loss function is defined as the sum of the data loss and physics loss. The data loss represents an error between observed data and the PINN predictions, whereas the physics loss enforces the derived governing equations by penalizing their residuals. Minimizing this loss simultaneously fits the observed dynamics and identifies the missing terms through the learned functions $g$ and $\theta$.
@@ 2.2 Variational principle and basic form of the governing equations
We employ the variational principle formulated by Fukagawa2010, Fukagawa2012, which can derive the Navier-Stokes equation including the inertia term. In the following, an overview of the derivation is described. The details of the derivation are provided in Appendix˜B.
We consider the dynamics of a viscous two-component fluid. Let $\rho$ and $\phi$ be the total mass density and the mass fraction of one component, respectively. The mass conservation can be given as
@@
where $\boldsymbol{u}$ and $\boldsymbol{j}$ are the velocity fields and the diffusive flux, respectively. The chemical potential is defined as the Gibbs energy per unit mass, and $\mu\left(\phi\right)$ is the difference of chemical potential between two components. In addition to the free energy due to the fluid motion, the free energy due to the two-component mixture is defined as $g(\phi)$, and we want to discover the form of $g\left(\phi\right)$ from the observed data. The total additional free energy of this system $G$ and the chemical potential $\mu$ are defined as
@@
$\rho T\left(\frac{\partial s}{\partial t}+\boldsymbol{u}\cdot\nabla s\right)-\theta+\nabla\cdot\left(\boldsymbol{J}_{q}+\mu\boldsymbol{j}\right)=0,$ (6)
where $\boldsymbol{J}$ is the heat flux, and $\theta$ is the dissipation function, which expresses how the frictional force converts the energy into heat. The total dissipation potential is defined as
@@
where $U\left(\boldsymbol{u},\rho,s\right)$ is the constraint and $\lambda$ is the Lagrangian multiplier. By solving the stationary condition of Eq.˜9, we can obtain the momentum equation. From the law of entropy, $\boldsymbol{j}\cdot\nabla\mu\leq 0$ must be satisfied. Assuming the linear phenomenological law, we can express $\boldsymbol{j}$ as
@@ 2.3 Architecture of GIMLET
Figure˜1 shows the architecture of the proposed method. The total framework consists of three neural networks: a PINN, FreeEnergyNet, and DissipationNet. The PINN, which is trainable through parameters $\boldsymbol{\lambda}$, learns the mapping of input coordinates $\{\boldsymbol{x},t\}$ to the solution $\boldsymbol{\psi}=\{u,v,w,p,\phi\}$. The FreeEnergyNet and DissipationNet learn the mapping of solutions to the free energy function $g$ and the dissipation function $\theta$, respectively. These neural networks are trainable through $\boldsymbol{\gamma}_{g}$ and $\boldsymbol{\gamma}_{\theta}$, respectively. The total loss function $\mathcal{J}$ is composed of the data loss $\mathcal{J}_{d}$, the physics loss $\mathcal{J}_{p}$, and the regularization terms $\mathcal{J}_{r}$. The data loss and physics loss are calculated using the measurement dataset $\mathcal{D}_{m}$ and residual points $\mathcal{D}_{f}$, respectively. By optimizing three neural networks simultaneously, the PINN is trained to infer the full solution field and the missing terms.
Fig. 1: Schematic architecture of the GIMLET for data-driven discovery of constitutive models in flow and transport PDEs.
Physics-informed neural networks (PINNs)
The PINNs methodology enables to impose the physical laws in the machine learning framework as soft constraints through a loss function. We use the PINN for the mapping of the solution field $\boldsymbol{\psi}$ from spatiotemporal coordinates $\{\boldsymbol{x},t\}$, and for taking spatial and time derivatives through AD. Among a lot of variations of PINNs, the original plain PINN [Raissi2019] is employed in this study.
FreeEnergyNet
The free energy function $g\left(\phi\right)$ is expressed by the FreeEnergyNet. Since only derivatives of $g\left(\phi\right)$ are required in the GEs, the FreeEnergyNet is composed of an Integrable Neural Network (INN) [Teichert2019]. The INN is a neural network designed such that its output can be precisely expressed as the gradient (or integral) of a certain scalar function. We consider the standard fully-connected neural network associated with multi-variable function $f\left(\boldsymbol{\xi}\right)=f\left(\xi_{1},\xi_{2},\ldots,\xi_{m}\right)$:
@@
where $\boldsymbol{W}^{\ell}_{y}$ and $\boldsymbol{W}^{\ell}_{w}$ are the weight matrices for hidden layers and passthrough layers, respectively. At the connection between the input and the first hidden layer, only the passthrough connection is considered, letting $\boldsymbol{W}_{y}^{0}=\boldsymbol{0}$. To guarantee the convexity with respect to $\boldsymbol{\xi}$, the weights $\boldsymbol{W}^{\ell}_{y}$ must be non-negative and the activation function $\sigma$ must be convex and non-decreasing.
Functional derivatives
@@
For these forms, the functional derivatives can be written by series expansions as
@@
For most cases of physical phenomena, truncation up to the second term is sufficient. All the derivatives in Eq.˜18 can be numerically evaluated by AD. The first term $\partial g/\partial\phi$ can be obtained without AD by use of the INN or the CINN, which give the derivatives directly. For the second term $\nabla\cdot\left(\partial g/\partial\left(\nabla\phi\right)\right)$, the gradient $\nabla\phi$ can be calculated by AD through the PINN. After $\partial g/\partial\left(\nabla\phi\right)$ is evaluated by the FreeEnergyNet, its divergence can be calculated by AD through both the PINN and FreeEnergyNet. One of the novelties in this study is that the evaluation of these functional derivatives is realized by the AD.
Loss functions and optimizers
@@
$\mathcal{D}_{m}=\left\{\boldsymbol{x}_{i}^{m},t_{i}^{m},\boldsymbol{\psi}_{i}^{m}\right\}_{i=1}^{N_{m}}.$ (21)
@@
$\mathcal{J}_{p}\left(\boldsymbol{\lambda},\boldsymbol{\gamma}_{g},\boldsymbol{\gamma}_{\theta};\mathcal{D}_{f}\right)$ $=\sum_{j}\frac{1}{N_{f}}\left\|\mathcal{F}_{j}\left[\boldsymbol{\psi}^{\lambda}\left(\boldsymbol{x}^{f},t^{f}\right),\boldsymbol{\gamma}_{g},\boldsymbol{\gamma}_{\theta}\right]\right\|_{2}^{2},$ (22b)
$\mathcal{J}_{d}\left(\boldsymbol{\lambda};\mathcal{D}_{m}\right)$ $=\frac{1}{N_{m}}\left\|\boldsymbol{\psi}^{\lambda}\left(\boldsymbol{x}^{m},t^{m}\right)-\boldsymbol{\psi}^{m}\right\|_{2}^{2},$ (22c)
@@
where $\mathcal{J}_{p}$ is the physics loss, which is defined as the mean-squared error (MSE) of the residual of GEs evaluated on residual points $\mathcal{D}_{f}$. $\mathcal{J}_{d}$ is the data loss, which is the MSE between observed data $\boldsymbol{\psi}^{m}$ and solution fields predicted by the PINN on data points $\{\boldsymbol{x}^{m},t^{m}\}$. $\mathcal{J}_{r}$ is the regularization loss, which consists of $\ell_{2}$ regularization $\mathcal{J}_{\ell_{2}}$, scalings for FreeEnergyNet $\mathcal{J}_{g}$ and DissipationNet $\mathcal{J}_{\theta}$, and pressure reference $\mathcal{J}_{h}$. These regularizations may not be used depending on the problems.

### v2

@@ Abstract
We develop a data-driven framework for discovering constitutive relations in models of fluid flow and scalar transport. Under the assumption that velocity and/or scalar fields are measured, our approach infers unknown closure terms in the governing equations as neural networks (gray-box discovery). The target to be discovered is the constitutive relations only, while the temporal derivative, convective transport terms, and pressure-gradient term in the governing equations are prescribed. The formulation is rooted in a variational principle from non-equilibrium thermodynamics, where the dynamics is defined by a free-energy functional and a dissipation functional. The unknown constitutive terms arise as functional derivatives of these functionals with respect to the state variables. To enable a flexible and structured model discovery, the free-energy and dissipation functionals are parameterized using neural networks, while their functional derivatives are obtained via automatic differentiation. This construction enforces thermodynamic consistency by design, guaranteeing monotonic decay of the total free energy and non-negative entropy production. The resulting method, termed GIMLET (Generalizable and Interpretable Model Learning through Embedded Thermodynamics), avoids reliance on a predefined library of candidate functions, unlike sparse regression or symbolic identification approaches. The learned models are generalizable in that functionals identified from one dataset can be transferred to distinct datasets governed by the same underlying equations. Moreover, the inferred free-energy and dissipation functions provide direct physical interpretability of the learned dynamics. The framework is demonstrated on several benchmark systems, including the viscous Burgers equation, the Kuramoto–Sivashinsky equation, and the incompressible Navier–Stokes equations for both Newtonian and non-Newtonian fluids.
@@ 1 Introduction
Neural operator style models aim to learn mappings between function spaces, such as the evolution of fields over time. Representative methods include Neural ODE [Chen2018], Hamiltonian Neural Networks [Greydanus2019], Fourier Neural Operator [Li2020], DeepONet [Lu2021], and many others [Lusch2018, Champion2019, Pfister2019, Yuan2019]. These approaches are largely black-box and often offer limited interpretability. In terms of generalizability, they typically require training on diverse data spanning many operating conditions, which can be time-consuming.
In contrast, Symbolic regression (SR) methods lie closer to the white box end of the spectrum and aim to recover explicit governing equations from data. Early successes in data-driven equation discovery were achieved using genetic programming-based symbolic regression, which represents candidate governing equations as expression trees and searches for parsimonious models consistent with data [Bongard2007, Schmidt2009]. SR comprises methods that discover mathematical expressions that best fit a given dataset [schmidt2009symbolic, stoutemyer2013can]. Among SR methods, genetic programming is one of the most widely used techniques, in which expression trees are iteratively modified using elementary mathematical operators to produce concise, data-consistent models. It expresses candidate models as combinations of elementary operators, variables, and constants, yielding explicit analytical forms. In genetic programming, candidate expressions are represented as tree structures, where internal nodes correspond to operators and leaf nodes represent variables or constants. Model discovery maintains a population of such trees and iteratively improves them through selection, mutation, and crossover based on their agreement with data. In recent work by [cranmer2023interpretable], both the performance and computational efficiency of symbolic regression were further enhanced through an island-based evolutionary framework. In this approach, populations were partitioned into multiple islands, each characterized by distinct operator sets and mutation strategies. Genetic algorithm operations, including mutation and crossover, were performed both within individual islands and across islands, promoting population diversity, improved exploration of the search space, and more effective convergence toward interpretable symbolic models.
@@
As reviewed so far, existing methods for physical model discovery typically fall short in at least one of three aspects: generalizability, interpretability, or library-free. We propose a new methodology that aims to achieve all three objectives. To this end, we leverage thermodynamics and a variational principle. Thermodynamics is among the most universal and transferable descriptions of physical systems. Enforcing thermodynamic consistency can lead to more robust constitutive laws, stable numerical schemes, and physically meaningful discovered models [Cueto2023]. Motivated by this idea, several machine learning frameworks have been proposed, including Themodynamics-based Artificial Neural Networks (TANNs) [Masi2021] and thermodynamically consistent physics-informed neural networks [Patel2022]. Motivated by the same goal of enforcing thermodynamic consistency in a general and systematic way, the General Equation for Non-Equilibrium Reversible Irreversible Coupling (GENERIC) framework was developed as a universal description of macroscopic dynamics. In GENERIC, the evolution equations are constructed from energy and entropy potentials, providing a unified structure that couples reversible and irreversible processes and can be viewed as a thermodynamically grounded extension of the Ginzburg–Landau equation [Grmela1997, Oettinger1997]. Based on the GENERIC concept, Zhang2022 proposed the GENERIC formalism-informed neural networks (GFINNs) for physical model discovery. However, the applicability of GFINN has so far been investigated primarily for ordinary differential equations, leaving extensions to PDE settings open.
According to the variational principle, also called the principle of least action path, realizable motions correspond to the stationary points of an action functional. Onsager’s variational principle [Onsager1931a, Onsager1931], which formulated the time evolution of the free energy and dissipation potentials, can lead to thermodynamically consistent governing equations for non-equilibrium dynamics [Van2020]. Building on this principle, Huang2022 proposed the Variational Onsager Neural Networks (VONNs), which aim to discover the free energy and dissipation potentials as neural networks. Although VONNs are a sophisticated and novel methodology, their application is limited to systems whose dynamics can be fully described by free energy and dissipation potentials. This restriction makes direct application to fluid flow challenging. For instance, applying Onsager’s variational principle to the Newtonian fluid in a straightforward way yields only a linearized form of the Navier–Stokes equation [Takaki2007]. In addition, the VONNs framework assumes that both the state variables and process variables are available as training data, which is often unrealistic for practical experimental observations.
@@ 2 Methodology
This section presents our thermodynamically consistent, library-free framework for discovering missing terms in continuum PDEs. We consider physical law discovery in continuum mechanics, where spatiotemporal field variables are evolving according to partial differential equations (PDEs). We assume that part of the governing equations is known and focus only on the missing terms (gray-box discovery). To represent these unknown contributions, an additional free energy function $g$ and dissipation function $\theta$ are introduced. Basic forms of governing equations (GEs) are derived through the variational principle. In the resulting GEs, the unknown terms are expressed as functional derivatives of $g$ and $\theta$. To learn $g$ and $\theta$ from data, we employ physics-informed neural networks (PINNs). Using PINNs, the functional derivatives for unknown functions can be calculated through automatic differentiation. The total loss function is defined as the sum of the data loss and physics loss. The data loss represents an error between observed data and the PINN predictions, whereas the physics loss enforces the derived governing equations by penalizing their residuals. Minimizing this loss simultaneously fits the observed dynamics and identifies the missing terms through the learned functions $g$ and $\theta$.
@@ 2.1 Variational principle and basic form of the governing equations
The variational principle formulated by Fukagawa2010, Fukagawa2012 is adopted, which can derive the Navier–Stokes equation including the inertia term. In the following, an overview of the derivation is described. The details of the derivation are provided in Appendix˜B. The dynamics of a viscous two-component fluid are considered. Let $\rho$ and $\phi$ be the total mass density and the mass fraction of one component, respectively. The mass conservation can be given as
@@
where $\boldsymbol{u}$ and $\boldsymbol{j}$ are the velocity fields and the diffusive flux, respectively. The chemical potential is defined as the Gibbs energy per unit mass, and $\mu\left(\phi\right)$ is the difference of chemical potential between two components. In addition to the free energy due to the fluid motion, the free energy due to the two-component mixture is defined as $g(\phi)$, and the form of $g\left(\phi\right)$ is inferred from the observed data. The total additional free energy of this system, $G$, and the chemical potential $\mu$ are defined as
@@
$\rho T\left(\frac{\partial s}{\partial t}+\boldsymbol{u}\cdot\nabla s\right)-\theta+\nabla\cdot\left(\boldsymbol{q}+\mu\boldsymbol{j}\right)=0,$ (6)
where $\boldsymbol{q}$ is the heat flux, and $\theta$ is the dissipation function, which expresses how the frictional force converts the energy into heat. The total dissipation potential is defined as
@@
where $U\left(\boldsymbol{u},\rho,s\right)$ is the constraint and $\lambda$ is the Lagrangian multiplier. By solving the stationary condition of Eq.˜9, the momentum equation is obtained. From the law of entropy, $\boldsymbol{j}\cdot\nabla\mu\leq 0$ must be satisfied. Assuming the linear phenomenological law, we can express $\boldsymbol{j}$ as
@@ 2.2 Architecture of GIMLET
Figure˜1 shows the architecture of the proposed method. The total framework consists of three neural networks: a PINN, FreeEnergyNet, and DissipationNet. The PINN, which is trainable through parameters $\boldsymbol{\lambda}$, learns the mapping of input coordinates $\{\boldsymbol{x},t\}$ to the solution $\boldsymbol{\psi}^{f}_{\boldsymbol{\lambda}}=\{u,v,w,p,\phi\}$. The FreeEnergyNet and DissipationNet learn the mapping of solutions to the free energy function $g$ and the dissipation function $\theta$, respectively. These neural networks are trainable through $\boldsymbol{\gamma}_{g}$ and $\boldsymbol{\gamma}_{\theta}$, respectively. The total loss function $\mathcal{J}$ is composed of the data loss $\mathcal{J}_{d}$, the physics loss $\mathcal{J}_{p}$, and the regularization terms $\mathcal{J}_{r}$. The data loss and physics loss are calculated using the measurement dataset $\mathcal{D}_{m}$ and residual points $\mathcal{D}_{f}$, respectively. By optimizing three neural networks simultaneously, the PINN is trained to infer the full solution field and the missing terms. The pressure field is assumed not to be measured, and the data loss is evaluated without pressure data. For the problem of fluid flow only, the scalar field $\phi$ and FreeEnergyNet are not used.
Fig. 1: Schematic architecture of GIMLET for data-driven discovery of constitutive models in flow and transport PDEs. The PINN predicts the solution fields $\psi^{f}_{\boldsymbol{\lambda}}$ from spatiotemporal inputs $\{\boldsymbol{x},t\}$, while FreeEnergyNet and DissipationNet parameterize the free-energy and dissipation functions. Automatic differentiation is used to evaluate functional derivatives, and training is performed by minimizing combined data and physics losses. The pressure field is assumed not to be measured, and the data loss is evaluated without pressure data. For the problem of fluid flow only, the scalar field $\phi$ and FreeEnergyNet are not used.
2.2.1 Physics-informed neural networks (PINNs)
The PINNs methodology enables imposing the physical laws in the machine learning framework as soft constraints through a loss function. We use the PINN for the mapping of the solution field $\boldsymbol{\psi}^{f}_{\lambda}$ from spatiotemporal coordinates $\{\boldsymbol{x},t\}$, and for taking spatial and time derivatives through AD. Among a lot of variations of PINNs, the original plain PINN [Raissi2019] is employed in this study.
2.2.2 FreeEnergyNet
The free energy function $g\left(\phi\right)$ is expressed by the FreeEnergyNet. Since only derivatives of $g\left(\phi\right)$ are required in the GEs, the FreeEnergyNet is composed of an Integrable Neural Network (INN) [Teichert2019]. The INN is a neural network designed such that its output can be precisely expressed as the gradient (or integral) of a certain scalar function. We consider the standard fully-connected neural network associated with a multi-variable function $f\left(\boldsymbol{\xi}\right)=f\left(\xi_{1},\xi_{2},\ldots,\xi_{m}\right)$:
@@ 2.2.3 DissipationNet
where $\boldsymbol{W}^{\ell}_{y}$ and $\boldsymbol{W}^{\ell}_{w}$ are the weight matrices for hidden layers and passthrough layers, respectively. At the connection between the input and the first hidden layer, only the passthrough connection is considered, letting $\boldsymbol{W}_{y}^{0}=\boldsymbol{0}$. To maintain the convexity with respect to $\boldsymbol{\xi}$, the weights $\boldsymbol{W}^{\ell}_{y}$ must be non-negative and the activation function $\sigma$ must be both convex and non-decreasing.
2.2.4 Functional derivatives
@@
For these forms, the functional derivatives can be written by using series expansions as
@@
For most cases of physical phenomena, truncation up to the second term is sufficient. All the derivatives in Eq.˜18 can be numerically evaluated by AD. The first term $\partial g/\partial\phi$ can be obtained without AD by using the INN or the CINN, which outputs the derivatives directly. For the second term $\nabla\cdot\left(\partial g/\partial\left(\nabla\phi\right)\right)$, the gradient $\nabla\phi$ can be calculated by AD through the PINN. After $\partial g/\partial\left(\nabla\phi\right)$ is evaluated by the FreeEnergyNet, its divergence can be calculated by AD through both the PINN and FreeEnergyNet. One of the novelties in this study is that the evaluation of these functional derivatives is realized by the AD.
2.2.5 Loss functions and optimizers
@@
$\mathcal{D}_{m}=\left\{\boldsymbol{x}_{i}^{m},t_{i}^{m},\boldsymbol{\psi}_{i}^{m}\right\}_{i=1}^{N_{m}},$ (21)
where $\boldsymbol{\psi}^{m}=\{u,v,w,\phi\}$ is the measured field data. Note that the pressure $p$ is not included in $\boldsymbol{\psi}^{m}$. In the practical observation of the flow field, such as Particle Image Velocimetry (PIV), obtaining the velocity and pressure at the same time seems to be extremely difficult. For this reason, only the velocity field is assumed to be measured, and pressure is not used in the evaluation of data loss. To be clear, the value of the pressure gradient is inferred by the PINN, which is substituted in the governing equation for the evaluation of physics loss.
@@
$\mathcal{J}_{p}\left(\boldsymbol{\lambda},\boldsymbol{\gamma}_{g},\boldsymbol{\gamma}_{\theta};\mathcal{D}_{f}\right)$ $=\sum_{j}\frac{1}{N_{f}}\left\|\mathcal{F}_{j}\left[\boldsymbol{\psi}^{f}_{\lambda}\left(\boldsymbol{x}^{f},t^{f}\right),\boldsymbol{\gamma}_{g},\boldsymbol{\gamma}_{\theta}\right]\right\|_{2}^{2},$ (22b)
$\mathcal{J}_{d}\left(\boldsymbol{\lambda};\mathcal{D}_{m}\right)$ $=\frac{1}{N_{m}}\left\|\boldsymbol{\psi}^{m}_{\lambda}\left(\boldsymbol{x}^{m},t^{m}\right)-\boldsymbol{\psi}^{m}\right\|_{2}^{2},$ (22c)
@@
where $\mathcal{J}_{p}$ is the physics loss, which is defined as the mean-squared error (MSE) of the residual of GEs evaluated on residual points $\mathcal{D}_{f}$. $\mathcal{J}_{d}$ is the data loss, which is the MSE between observed data $\boldsymbol{\psi}^{m}$ and solution fields $\boldsymbol{\psi}^{m}_{\lambda}$ predicted by the PINN on data points $\{\boldsymbol{x}^{m},t^{m}\}$. $\mathcal{J}_{r}$ is the regularization loss, which consists of $\ell_{2}$ regularization $\mathcal{J}_{\ell_{2}}$, scalings for FreeEnergyNet $\mathcal{J}_{g}$ and DissipationNet $\mathcal{J}_{\theta}$, and pressure reference $\mathcal{J}_{h}$. $\beta_{\ell_{2}},\beta_{g},\beta_{\theta},\beta_{h}$ are the weighting factor for the corresponding regularizations. These regularizations may not be used depending on the problems.